Find the area of the triangle with the base 1 4/5 yards and height 5/8 yard
Answer by Alan3354(69443)

Find the area of the triangle with the base 1 4/5 yards and height 5/8 yard
---------------
Area = b*h/2 = 1.8*0.625 sq yds